OpenStudy (anonymous):

i need help solving this pleasee 4^2x = 32

OpenStudy (yoloshroom):

So is this your question? \[4^{2x}=32\]

OpenStudy (anonymous):

4^2= 16, because 4*4 is 16. Now, you divide both sides by 16. \[\frac{ 16x }{ 16 }=\frac{ 32 }{ 16 }\] 16x divided by 16 cancels out. NOw, what would 32 divided by 16 be?

Nnesha (nnesha):

it's 4 to the (2x) power :=)

OpenStudy (anonymous):

XD I JUST noticed that!

OpenStudy (anonymous):

My bad :p

Nnesha (nnesha):

it's okay and selly somebody asked a question you should answer it.

OpenStudy (campbell_st):

write both sides of the equation as powers of 2. \[4^{2x} = 32 ~~~or~~~~(2^2)^{2x} = 2^5\] so after multiplying the powers you get \[2^{4x} = 2^5\] equate the powers so 4x = 5 and just solve for x

OpenStudy (anonymous):

yes thats my question sorry guys i had to take a moment to change my babies diapers ! anyways but oh ok i see whats your saying. so my answer is x=0.8

OpenStudy (anonymous):

or just x=5/4 actually

OpenStudy (campbell_st):

well done